OpenStudy (anonymous):

A purse contains 21 coins consisting of nickels and dimes. Find the number of coins of each kind if the total value is $1.65. If 5n represents the value of the nickels, then which of the following expressions represents the value of the dimes? A.10n B.16n C.10(21 - n)

OpenStudy (anonymous):

nickles + dimes = 21 coins n + d = 21 (eq 1) let n be the number of nickels let d be the number of dimes 0.05*n + 0.10*d= 1.65 (eq 2) 2 equations, 2 unknowns, so, substitute equation (1) into equation (2) solve for either n or d with the new eqation made

OpenStudy (anonymous):

I wrote a Python script and spun through the possibilities: 9 nickels, 12 dimes.

OpenStudy (anonymous):

*intresting...